On complete combustion, 0.246 g of an organic compound gave 0.198 g of CO2 and 0.1014 g of H2O. What is the ratio of carbon and hydrogen atoms in the compound?


1
Expert's answer
2022-04-08T16:00:05-0400

Mass fraction of C in CO₂ = 12/(12 + 16×2) = 12/44

Mass fraction of H in H₂O = 1×2/(1×2 + 16) = 2/18

Mass of C in the compound = Mass of C in CO₂ = (0.246 g) × (12/44) = 0.0671 g

Mass of H in the compound = Mass of H in H₂O = (0.1014 g) × (2/18) = 0.0113 g

In the compound, the ratio of carbon and hydrogen atoms

= Mole ratio C : H

= (0.0671/12) : (0.0113/1)

= 0.00559 : 0.0113 …… (0.00559 is the smallest number)

= (0.00559/0.00559) : (0.0113/0.00559)

= 1 : 2